Can you show me the steps to solve 5 over 3x + 3 over x = 1? Thanks!
Answer by algebrahouse.com(1659)

5/3x + 3/x = 1
3x[5/3x + 3/x = 1] {multiplied entire equation by least common denominator, 3x}
3x(5/3x) + 3x(3/x) = 3x(1) {multiplied each term by 3x}
5 + 9 = 3x {cancelled and multiplied what's left}
3x = 14 {added 5 and 9}
x = 14/3 {divided both sides by 3}

For more help from me, visit: www.algebrahouse.com